# Homological stratification and descent

### v1

@@ 5. Comparison with the Balmer–Favi support
In this section we will compare the homological support with the Balmer–Favi support and, in particular, relate the h-detection and h-LGP properties with the usual detection and LGP properties. We assume some familiarity with [BHS23b].
@@
For any Thomason subset $Y\subseteq\operatorname{Spc}(\mathscr{T}^{c})$, [BHS23a, Lemmma 3.8] establishes that $\operatorname{Supp}^{h}(e_{Y})=\pi^{-1}(Y)$ and $\operatorname{Supp}^{h}(f_{Y})=\pi^{-1}(Y^{c})$. Moreover, we have $\operatorname{Supp}^{h}(e_{Y})=\operatorname{Supp}^{n}(e_{Y})$ and $\operatorname{Supp}^{h}(f_{Y})=\operatorname{Supp}^{n}(f_{Y})$ since $e_{Y}$ is a (weak) coring and $f_{Y}$ is a (weak) ring (1.7). Finally, we claim that if $a,b\in\mathscr{T}$ both have the property that their naive and genuine h-supports coincide, then so does their tensor product $a\otimes b$. Indeed,
@@
For the objects $a$ under consideration, the naive and genuine h-support coincide by 5.2 and 5.3, so we actually have
@@
Otherwise, $\mathsf{hom}(a,\mathsf{hom}(E_{\mathscr{B}},t))=0$ so that $\mathsf{hom}(e_{\operatorname{supp}(a)},\mathsf{hom}(E_{\mathscr{B}},t))=0$. Hence $\mathsf{hom}(E_{\mathscr{B}},t)\simeq\mathsf{hom}(f_{\operatorname{supp}(a)},\mathsf{hom}(E_{\mathscr{B}},t))$. But $\mathsf{hom}(E_{\mathscr{B}},t)\neq 0$ by hypothesis. It follows that $E_{\mathscr{B}}\otimes f_{\operatorname{supp}(a)}\neq 0$. That is, $\mathscr{B}\in\operatorname{Supp}^{n}(f_{\operatorname{supp}(a)})=\pi^{-1}(\operatorname{supp}(a)^{c})=\pi^{-1}(\operatorname{supp}(a))^{c}=\operatorname{Supp}^{h}(a)^{c}$ by 5.3, which contradicts the hypothesis. This proves the claim when $a$ is compact.
@@
We proved in [BHS23a, Proposition 3.10] that the inclusion $\pi(\operatorname{Supp}^{h}(t))\subseteq\operatorname{Supp}(t)$ always holds. On the other hand, consider the purported inclusion
@@
by 5.5 and 5.3. Thus, $\mathscr{P}\not\in\pi(\operatorname{Cosupp}^{h}(t))$.
@@
by 5.5 and 5.3 so that $\mathscr{P}\in\pi(\operatorname{Cosupp}^{h}(t))$. ∎
@@
This immediately follows from 5.6. ∎
@@
The detection property does not in general imply the h-detection property and the codetection property does not in general imply the h-codetection property. Indeed, as explained in [BHS23a, Example 5.5], Neeman [Nee00] provides an example of a non-noetherian commutative ring $R$ whose spectrum $\operatorname{Spec}(R)$ is a single point with the property that $\operatorname{D}(R)$ contains a nontrivial tensor-nilpotent object $I\in\operatorname{D}(R)$; cf. 3.14. The tensor product property forces ${\operatorname{Supp}^{h}(I)=\varnothing}$. Thus $\operatorname{D}(R)$ does not have the h-detection property and hence it does not have the h-LGP property or h-codection property (2.5), either. However, by [BHS23b, Theorem 3.22] it does have the LGP (which is equivalent to the codetection property by [BCHS23, Theorem 6.4]) since the spectrum $\operatorname{Spc}(\operatorname{D}(R)^{c})=\operatorname{Spec}(R)=*$ is a noetherian space.
5.9.
@@
The $\subseteq$ inclusion always holds (5.6), so it suffices to establish $\pi(\operatorname{Supp}^{h}(w))\supseteq\operatorname{Supp}(w)$. Let $\mathscr{P}\in\operatorname{Supp}(w)$. We have $\{\mathscr{P}\}=\operatorname{supp}(x)\cap\operatorname{gen}(\mathscr{P})$ for some $x\in\mathscr{T}^{c}$ by [BHS23b, Remark 2.8]. Moreover, replacing $x$ by $x\otimes x^{\vee}$ we may assume that $x$ is a (weak) ring. Then $w\otimes g_{\mathscr{P}}\neq 0$ implies that $w\otimes x\otimes f_{\operatorname{gen}(\mathscr{P})^{c}}\neq 0$ is a nonzero weak ring. The homological support has the detection property for weak rings by [Bal20a, Theorem 4.7]. Hence
@@
by the tensor product property (1.1), 5.2 and 5.3. Therefore $\mathscr{P}\in\pi(\operatorname{Supp}^{h}(w))$, as desired. ∎
@@ 6. Base change for homological support
For any homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{S}^{c})$, $E_{\varphi^{h}(\mathscr{B})}\in\mathscr{T}$ is a direct summand of ${f_{*}(E_{\mathscr{B}})\in\mathscr{T}}$.
@@
The corollary shows that the Avrunin–Scott identity holds unconditionally whenever $\mathscr{T}$ has the h-detection property. This contrasts significantly with the situation for the Balmer–Favi support; cf. [BCHS23, Corollary 14.19] and [BCH+24b, Corollary 12.8].
@@
The equivalence of $(b)$ and $(c)$ is [BCHS24, Theorem 1.8] and $(a)$ certainly implies $(b)$. The implication $(c)\Rightarrow(a)$ follows from 6.4. ∎
@@
    $\varphi^{h}(\operatorname{Supp}^{h}(w))\subseteq\operatorname{Supp}^{h}(f_{*}(w))$ for any weak ring $w\in\mathscr{S}$.
@@
For the first statement, let $\mathscr{B}\in\varphi^{h}(\operatorname{Supp}^{h}(t))$. Using 6.4, we then have
$\operatorname{Supp}^{h}(f^{*}(E_{\mathscr{B}})\otimes w)=(\varphi^{h})^{-1}(\{\mathscr{B}\})\cap\operatorname{Supp}^{h}(w)\neq\varnothing.$
Thus $f^{*}(E_{\mathscr{B}})\otimes w\neq 0$. This implies $f_{*}(f^{*}(E_{\mathscr{B}})\otimes w)\neq 0$ since $w$ is a weak ring (see [BCHS23, Remark 13.12]). Hence $E_{\mathscr{B}}\otimes f_{*}(w)\neq 0$. This implies $\mathscr{B}\in\operatorname{Supp}^{h}(f_{*}(w))$ since $f_{*}(w)$ is a weak ring.
@@
Recall from [Bal20b, Remark 3.4] and [BHS23a, Example 2.8] that a basis of closed sets for the topology on $\operatorname{Spc}^{h}(\mathscr{S}^{c})$ is given by the $\operatorname{Supp}^{h}(x)$ for $x\in\mathscr{S}^{c}$. Since $f^{*}$ is a finite localization, $x\oplus\Sigma x=f^{*}(c)$ for some compact $c\in\mathscr{T}^{c}$ and (replacing $c$ with $c\otimes c^{\vee}$) we can assume $c$ is a (weak) ring. Then by 6.4 we have
@@
To describe the image of $\varphi^{h}$, note that the object $f_{*}(\mathbb{1}_{\mathscr{S}})$ is the right idempotent of the finite localization so $\operatorname{Supp}^{h}(f_{*}(\mathbb{1}_{\mathscr{S}}))=\pi^{-1}(\operatorname{im}(\varphi))$ by 5.3 and $\operatorname{Supp}^{h}(f_{*}(\mathbb{1}_{\mathscr{S}}))=\operatorname{im}(\varphi^{h})$ by [Bal20a, Theorem 5.12]. ∎
@@ 7. Weakly descendable functors and descent
For a geometric family $(f_{i}^{*})_{i\in I}$, the following conditions are equivalent:
@@
    $(f_{i}^{!})_{i\in I}$ is jointly conservative.
@@
Write $\mathscr{L}=\operatorname{Locid}\langle(f_{i})_{*}\mathbb{1}\mid i\in I\rangle$ and consider its right orthogonal $\mathscr{L}^{\perp}$ in $\mathscr{T}$. Note that
@@
The first equivalence follows from [BCHS23, Definition 2.9], the second equivalence follows from the internal adjunction [BDS16, (2.18)], and the final equivalence follows from the observation that a left adjoint is conservative on the essential image of its right adjoint, by the unit-counit identity. ∎
@@
$(a)\Rightarrow(b)$: This follows from 7.5 and the proof of [BCHS23, Proposition 13.21]. $(b)\Rightarrow(c)$: This is immediate from the definitions. $(c)\Leftrightarrow(d)$: This is [BCHS24, Theorem 1.8]. $(d)\Rightarrow(e)$: This holds because $\pi$ is surjective. $(e)\Rightarrow(a)$: Since $\varphi$ is surjective, we have
$\operatorname{Spc}(\mathscr{T}^{c})=\operatorname{im}\varphi=\bigcup_{i\in I}\operatorname{im}\varphi_{i}=\bigcup_{i\in I}\operatorname{Supp}((f_{i})_{*}(\mathbb{1})),$
where the final equality uses [BCHS23, Corollary 13.15]. If $f_{i}^{*}$ is weakly closed for all $i\in I$ then $\mathscr{T}^{c}=\operatorname{Thickid}\langle(f_{i})_{*}\mathbb{1}\mid i\in I\rangle$ by the classification of thick ideals. ∎
@@
In general, $(f_{i}^{*})_{i\in I}$ being jointly conservative does not imply that the family is weakly descendable. Indeed, let $k$ be a field and $I$ an infinite set. By [Ste14], the projection functors
@@
The next result is a minor variation on 7.8 which improves on [BCHS23, Corollary 14.24] by extending the latter to families of geometric functors and weakening the tt-stratification assumption to h-stratification.
@@
The implications $(a)\Rightarrow(b)\Rightarrow(c)\Rightarrow(d)$ are part of 7.8, so it remains to show $(d)\Rightarrow(a)$. Indeed, if $\varphi^{h}$ is surjective, then
@@
In light of the above corollary and its analogue [BCHS23, Corollary 14.24], we regard weak descendability as the appropriate condition for studying descent properties in rigidly-compactly generated tt-geometry.
@@
The converse of 7.15 does not hold. For example, the homological local-to-global principle does not hold for the derived category of the truncated polynomial ring in 3.14 (see 3.15), but the family of localizations is weakly descendable for the trivial reason that the category is local.
@@
If $\mathscr{T}$ has the h-detection property then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle(f_{\mathscr{B}})_{*}\mathbb{1}\rangle$ for every $\mathscr{B}$ by 3.5, which implies the desired equivalence. Certainly h-LGP implies h-detection. To finish the proof, we claim that $(a)$ also implies h-detection. Indeed, since tt-fields are h-stratified (3.4), we have $(\varphi^{h}_{\mathscr{B}})^{-1}(\operatorname{Supp}^{h}(t))=\operatorname{Supp}^{h}(f_{\mathscr{B}}^{*}(t))$ for every $\mathscr{B}$ and every object $t\in\mathscr{T}$ by 6.8. Therefore, the h-detection property is equivalent to the family $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$ being jointly conservative. The claim thus follows from 7.8. ∎
@@
3.1$(b)$ applied to $\mathscr{S}_{i}$ then shows that
@@
After applying $(f_{i})_{*}$ to both sides and employing the projection formula, the last displayed equality is equivalent to
$\operatorname{Locid}\langle t\otimes(f_{i})_{*}(\mathbb{1})\rangle=\operatorname{Locid}\langle E_{\mathscr{B}}\otimes(f_{i})_{*}(\mathbb{1})\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle.$
Varying now over $i\in I$ and keeping in mind that (7.1) holds by assumption, we get $\operatorname{Locid}\langle t\rangle=\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$, as desired. ∎
7.22.
We also include an alternative proof of 7.21 based on cosupport.
Alternative proof of 7.21.
We will verify condition (c) of 4.5. Note that the collection of maps $(\varphi_{i}^{h})$ is jointly surjective by 7.8. Since $(f_{i}^{*})$ is weakly descendable and $\mathscr{S}_{i}$ is h-stratified for all $i$, for any objects $t_{1},t_{2}\in\mathscr{T}$ we have equivalences:

### v2

@@ 5. Comparison with the Balmer–Favi support
In this section we will compare the homological support with the Balmer–Favi support and, in particular, relate the h-detection and h-LGP properties with the usual detection and LGP properties. We assume some familiarity with [BHS23d].
@@
Remark 5.3.
A subset $W\subseteq\operatorname{Spc}(\mathscr{T}^{c})$ is weakly visible if it can be written as the intersection of a Thomason subset and the complement of a Thomason subset: $W=Y_{1}\cap Y_{2}^{c}$. In this case, the object $g_{W}\coloneqq e_{Y_{1}}\otimes f_{Y_{2}}\in\mathscr{T}$ only depends on $W$ up to isomorphism. A point $\mathscr{P}\in\operatorname{Spc}(\mathscr{T}^{c})$ is said to be weakly visible if the singleton subset $\{\mathscr{P}\}$ is weakly visible, and in this case we write $g_{\mathscr{P}}\coloneqq g_{\{\mathscr{P}\}}$. Finally, the space $\operatorname{Spc}(\mathscr{T}^{c})$ is weakly noetherian if every point is weakly visible. See [BCH+23b, Section 4] for further details.
@@
For any Thomason subset $Y\subseteq\operatorname{Spc}(\mathscr{T}^{c})$, [BHS23c, Lemmma 3.8] establishes that $\operatorname{Supp}^{h}(e_{Y})=\pi^{-1}(Y)$ and $\operatorname{Supp}^{h}(f_{Y})=\pi^{-1}(Y^{c})$. Moreover, we have $\operatorname{Supp}^{h}(e_{Y})=\operatorname{Supp}^{n}(e_{Y})$ and $\operatorname{Supp}^{h}(f_{Y})=\operatorname{Supp}^{n}(f_{Y})$ since $e_{Y}$ is a (weak) coring and $f_{Y}$ is a (weak) ring (1.7). Finally, we claim that if $a,b\in\mathscr{T}$ both have the property that their naive and genuine h-supports coincide, then so does their tensor product $a\otimes b$. Indeed,
@@
For the objects $a$ under consideration, the naive and genuine h-support coincide by 5.2 and 5.4, so we actually have
@@
Otherwise, $\mathsf{hom}(a,\mathsf{hom}(E_{\mathscr{B}},t))=0$ so that $\mathsf{hom}(e_{\operatorname{supp}(a)},\mathsf{hom}(E_{\mathscr{B}},t))=0$. Hence $\mathsf{hom}(E_{\mathscr{B}},t)\simeq\mathsf{hom}(f_{\operatorname{supp}(a)},\mathsf{hom}(E_{\mathscr{B}},t))$. But $\mathsf{hom}(E_{\mathscr{B}},t)\neq 0$ by hypothesis. It follows that $E_{\mathscr{B}}\otimes f_{\operatorname{supp}(a)}\neq 0$. That is, $\mathscr{B}\in\operatorname{Supp}^{n}(f_{\operatorname{supp}(a)})=\pi^{-1}(\operatorname{supp}(a)^{c})=\pi^{-1}(\operatorname{supp}(a))^{c}=\operatorname{Supp}^{h}(a)^{c}$ by 5.4, which contradicts the hypothesis. This proves the claim when $a$ is compact.
@@
We proved in [BHS23c, Proposition 3.10] that the inclusion $\pi(\operatorname{Supp}^{h}(t))\subseteq\operatorname{Supp}(t)$ always holds. On the other hand, consider the purported inclusion
@@
by 5.6 and 5.4. Thus, $\mathscr{P}\not\in\pi(\operatorname{Cosupp}^{h}(t))$.
@@
by 5.6 and 5.4 so that $\mathscr{P}\in\pi(\operatorname{Cosupp}^{h}(t))$. ∎
@@
This immediately follows from 5.7. ∎
@@
The detection property does not in general imply the h-detection property and the codetection property does not in general imply the h-codetection property. Indeed, as explained in [BHS23c, Example 5.5], Neeman [NEE00] provides an example of a non-noetherian commutative ring $R$ whose spectrum $\operatorname{Spec}(R)$ is a single point with the property that $\operatorname{D}(R)$ contains a nontrivial tensor-nilpotent object $I\in\operatorname{D}(R)$; cf. 3.14. The tensor product property forces $\operatorname{Supp}^{h}(I)=\varnothing$}. Thus $\operatorname{D}(R)$ does not have the h-detection property and hence it does not have the h-LGP property or h-codection property (2.5), either. However, by [BHS23d, Theorem 3.22] it does have the LGP (which is equivalent to the codetection property by [BCH+23b, Theorem 6.4]) since the spectrum $\operatorname{Spc}(\operatorname{D}(R)^{c})=\operatorname{Spec}(R)=*$ is a noetherian space.
5.10.
@@
The $\subseteq$ inclusion always holds (5.7), so it suffices to establish $\pi(\operatorname{Supp}^{h}(w))\supseteq\operatorname{Supp}(w)$. Let $\mathscr{P}\in\operatorname{Supp}(w)$. We have $\{\mathscr{P}\}=\operatorname{supp}(x)\cap\operatorname{gen}(\mathscr{P})$ for some $x\in\mathscr{T}^{c}$ by [BHS23d, Remark 2.8]. Moreover, replacing $x$ by $x\otimes x^{\vee}$ we may assume that $x$ is a (weak) ring. Then $w\otimes g_{\mathscr{P}}\neq 0$ implies that $w\otimes x\otimes f_{\operatorname{gen}(\mathscr{P})^{c}}\neq 0$ is a nonzero weak ring. The homological support has the detection property for weak rings by [BAL20a, Theorem 4.7]. Hence
@@
by the tensor product property (1.1), 5.2 and 5.4. Therefore $\mathscr{P}\in\pi(\operatorname{Supp}^{h}(w))$, as desired. ∎
@@ 6. Base change for homological support
Let $f^{*}:\mathscr{T}\to\mathscr{S}$ be a geometric functor. For any homological prime $\mathscr{B}\in\operatorname{Spc}^{h}(\mathscr{S}^{c})$, $E_{\varphi^{h}(\mathscr{B})}\in\mathscr{T}$ is a direct summand of $f_{*}(E_{\mathscr{B}})\in\mathscr{T}$}.
@@
The above corollary (6.4) shows that the Avrunin–Scott identity holds unconditionally whenever $\mathscr{T}$ has the h-detection property. This contrasts significantly with the situation for the Balmer–Favi support; cf. [BCH+23b, Corollary 14.19] and [BCH+24a, Corollary 12.8].
@@
The equivalence of $(b)$ and $(c)$ is [BCH+24b, Theorem 1.8] and $(a)$ certainly implies $(b)$. The implication $(c)\Rightarrow(a)$ follows from 6.4. ∎
@@
    $\varphi^{h}(\operatorname{Supp}^{h}(R))=\operatorname{Supp}^{h}(f_{*}(R))$ for any weak ring $R\in\mathscr{S}$.
@@
For the first statement, observe that
$\mathscr{B}\in\varphi^{h}(\operatorname{Supp}^{h}(R))$ $\Longleftrightarrow(\varphi^{h})^{-1}(\{\mathscr{B}\})\cap\operatorname{Supp}^{h}(R)\neq\varnothing$
$\Longleftrightarrow\operatorname{Supp}^{h}(f^{*}(E_{\mathscr{B}})\otimes R)\neq\varnothing$ (6.4)
$\Longleftrightarrow f^{*}(E_{\mathscr{B}})\otimes R\neq 0$ (1.8)
$\Longleftrightarrow f_{*}(f^{*}(E_{\mathscr{B}})\otimes R)\neq 0$ ([BCH+23b, Remark 13.12])
$\Longleftrightarrow E_{\mathscr{B}}\otimes f_{*}(R)\neq 0$
$\Longleftrightarrow\mathscr{B}\in\operatorname{Supp}^{h}(f_{*}(R)).$
@@
Recall from [BAL20b, Remark 3.4] and [BHS23c, Example 2.8] that a basis of closed sets for the topology on $\operatorname{Spc}^{h}(\mathscr{S}^{c})$ is given by the $\operatorname{Supp}^{h}(x)$ for $x\in\mathscr{S}^{c}$. Since $f^{*}$ is a finite localization, $x\oplus\Sigma x=f^{*}(c)$ for some compact $c\in\mathscr{T}^{c}$ and (replacing $c$ with $c\otimes c^{\vee}$) we can assume $c$ is a (weak) ring. Then by 6.4 we have
@@
To describe the image of $\varphi^{h}$, note that the object $f_{*}(\mathbb{1}_{\mathscr{S}})$ is the right idempotent of the finite localization so $\operatorname{Supp}^{h}(f_{*}(\mathbb{1}_{\mathscr{S}}))=\pi^{-1}(\operatorname{im}(\varphi))$ by 5.4 and $\operatorname{Supp}^{h}(f_{*}(\mathbb{1}_{\mathscr{S}}))=\operatorname{im}(\varphi^{h})$ by [BAL20a, Theorem 5.12]. ∎
@@ 7. Weakly descendable functors and descent
For a geometric family $(f_{i}^{*})_{i\in I}$, the following are equivalent:
@@
    $(f_{i}^{!})_{i\in I}$ is jointly conservative;
  3. (c)
    for any $t_{1},t_{2}\in\mathscr{T}$, we have
    $t_{1}\in\operatorname{Locid}\langle t_{2}\rangle\iff\forall i\in I\colon f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle f_{i}^{*}(t_{2})\rangle.$
@@
We first prove $(a)\Leftrightarrow(b)$. Write $\mathscr{L}=\operatorname{Locid}\langle(f_{i})_{*}\mathbb{1}\mid i\in I\rangle$ and consider its right orthogonal $\mathscr{L}^{\perp}$ in $\mathscr{T}$. Note that
@@
The first equivalence follows from [BCH+23b, Definition 2.9], the second equivalence follows from the internal adjunction [BDS16, (2.18)], and the final equivalence follows from the observation that a left adjoint is conservative on the essential image of its right adjoint, by the unit-counit identity.
$(a)\Rightarrow(c)$: If $t_{1}\in\operatorname{Locid}\langle t_{2}\rangle$ then $f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle f_{i}^{*}(t_{2})\rangle$ for all $i\in I$, since $f_{i}^{*}$ are tt-functors, so it remains to prove the reverse implication in $(c)$. Tensoring (7.1) with $t_{1}$ and using the projection formula gives
$t_{1}\in\operatorname{Locid}\langle t_{1}\otimes(f_{i})_{*}f_{i}^{*}(\mathbb{1})\mid i\in I\rangle=\operatorname{Locid}\langle(f_{i})_{*}f_{i}^{*}(t_{1})\mid i\in I\rangle.$
Assuming $f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle f_{i}^{*}(t_{2})\rangle$ and using the projection formula again, we deduce $(f_{i})_{*}f_{i}^{*}(t_{1})\in\operatorname{Locid}\langle(f_{i})_{*}f_{i}^{*}(t_{2})\rangle$. Now varying over $i\in I$ yields
$t_{1}\in\operatorname{Locid}\langle(f_{i})_{*}f_{i}^{*}(t_{1})\mid i\in I\rangle$ $\subseteq\operatorname{Locid}\langle(f_{i})_{*}f_{i}^{*}(t_{2})\mid i\in I\rangle$
$=\operatorname{Locid}\langle t_{2}\otimes(f_{i})_{*}f_{i}^{*}(\mathbb{1})\mid i\in I\rangle$
$\subseteq\operatorname{Locid}\langle t_{2}\rangle,$
as claimed.
$(c)\Rightarrow(a)$: By the triangle identity for the $f_{i}^{*}\dashv(f_{i})_{*}$ adjunction, $f_{i}^{*}(\mathbb{1})$ is a retract of $f_{i}^{*}(f_{i})_{*}f_{i}^{*}(\mathbb{1})\simeq f_{i}^{*}(f_{i})_{*}(\mathbb{1})$, hence $f_{i}^{*}(\mathbb{1})\in\operatorname{Locid}\langle f_{i}^{*}(f_{i})_{*}(\mathbb{1}))\rangle$ for all $i\in I$. Therefore, condition $(c)$ implies $\mathbb{1}\in\operatorname{Locid}\langle(f_{i})_{*}(\mathbb{1})\mid i\in I\rangle$, so $(f_{i}^{*})_{i\in I}$ is weakly descendable. ∎
@@
Remark 7.7.
Condition $(c)$ of 7.5 has been used in [BBI+23a] to establish a fiberwise criterion for stratification. Its $t$-geometric incarnation is studied further in work of Juan Omar Gómez [GÓM25].
@@
$(a)\Rightarrow(b)$: This follows from 7.5 and the proof of [BCH+23b, Proposition 13.21]. $(b)\Rightarrow(c)$: This is immediate from the definitions. $(c)\Leftrightarrow(d)$: This is [BCH+24b, Theorem 1.8]. $(d)\Rightarrow(e)$: This holds because $\pi$ is surjective. $(e)\Rightarrow(a)$: Since $\varphi$ is surjective, we have
$\operatorname{Spc}(\mathscr{T}^{c})=\operatorname{im}\varphi=\bigcup_{i\in I}\operatorname{im}\varphi_{i}.$
If the $f_{i}^{*}$ are weakly closed then, by definition, each $(f_{i})_{*}(\mathbb{1})$ is compact. We then have $\operatorname{im}\varphi_{i}\subseteq\operatorname{supp}((f_{i})_{*}(\mathbb{1}))$ by the argument given at the beginning of the proof of [BAL18, Theorem 1.7]. Hence
$\operatorname{Spc}(\mathscr{T}^{c})=\bigcup_{i\in I}\operatorname{im}\varphi_{i}\subseteq\bigcup_{i\in I}\operatorname{supp}((f_{i})_{*}(\mathbb{1}))$
so that $\mathscr{T}^{c}=\operatorname{Thickid}\langle(f_{i})_{*}\mathbb{1}\mid i\in I\rangle$ by the classification of thick ideals. ∎
@@
In general, $(f_{i}^{*})_{i\in I}$ being jointly conservative does not imply that the family is weakly descendable. Indeed, let $k$ be a field and II an infinite set. By [STE14], the projection functors
@@
The next result is a minor variation on 7.9 which improves on [BCH+23b, Corollary 14.24] by extending the latter to families of geometric functors and weakening the $t$-stratification assumption to h-stratification.
@@
The implications $(a)\Rightarrow(b)\Rightarrow(c)\Rightarrow(d)$ are part of 7.9, so it remains to show $(d)\Rightarrow(a)$. Indeed, if $\varphi^{h}$ is surjective, then
@@
In light of the above corollary and its analogue [BCH+23b, Corollary 14.24], we regard weak descendability as the appropriate condition for studying descent properties in rigidly-compactly generated tt-geometry.
@@
The converse of 7.16 does not hold. For example, the homological local-to-global principle does not hold for the derived category of the truncated polynomial ring in 3.14 (see 3.15), but the family of localizations is weakly descendable for the trivial reason that the category is local.
@@
If $\mathscr{T}$ has the h-detection property then $\operatorname{Locid}\langle E_{\mathscr{B}}\rangle=\operatorname{Locid}\langle(f_{\mathscr{B}})_{*}\mathbb{1}\rangle$ for every $\mathscr{B}$ by 3.5, which implies the desired equivalence. Certainly h-LGP implies h-detection. To finish the proof, we claim that $(a)$ also implies h-detection. Indeed, since tt-fields are h-stratified (3.4), we have $(\varphi^{h}_{\mathscr{B}})^{-1}(\operatorname{Supp}^{h}(t))=\operatorname{Supp}^{h}(f_{\mathscr{B}}^{*}(t))$ for every $\mathscr{B}$ and every object $t\in\mathscr{T}$ by 6.8. Therefore, the h-detection property is equivalent to the family $\{f_{\mathscr{B}}^{*}\colon\mathscr{T}\to\mathscr{F}_{\mathscr{B}}\}_{\mathscr{B}}$ being jointly conservative. The claim thus follows from 7.9. ∎
@@
3.1(b)(b) applied to $\mathscr{S}_{i}$ then shows that
@@
Keeping in mind that (7.1) holds by assumption, 7.5 implies that $\operatorname{Locid}\langle t\rangle=\operatorname{Locid}\langle E_{\mathscr{B}}\mid\mathscr{B}\in\operatorname{Supp}^{h}(t)\rangle$, as desired. ∎
7.23.
We also include an alternative proof of 7.22 based on cosupport.
Alternative proof of 7.22.
We will verify condition (c) of 4.5. Note that the collection of maps $(\varphi_{i}^{h})$ is jointly surjective by 7.9. Since $(f_{i}^{*})$ is weakly descendable and $\mathscr{S}_{i}$ is h-stratified for all $i$, for any objects $t_{1},t_{2}\in\mathscr{T}$ we have equivalences: